If nominal GDP increased by 11 percent and real GDP increased by only 5 percent, the price level must have increased by tentatively around 6 percent. Why is this so?

By using GDP deflator, the price level has increased by (11/5) X 100 = 220. Assuming everything is kept constant ceteris paribus. I don’t see why the concept of GDP deflator is not being able to be used here as we are speaking about price level.
1
Expert's answer
2018-07-04T09:00:09-0400
Calculations can be done in two ways:
Method 1 (approximate).
The inflation rate will be 11-5 = 6 (%).
Method 2 (accurate).
Denote the old real GDP as x, and the old nominal GDP as y.
Then the old value of the deflator is y / x, the new real GDP is equal to 1.05x, the new nominal GDP is equal to 1.11y, the new value of the deflator is equal to 1.11y / 1.05x = 1.0571 (y / x).
So, the inflation rate will be 5.71% - rounded to 6%.